f(x) = x2 + 3; g(x)= square root x-2
Find f(g(x)).

Help me solve this please
1
Expert's answer
2014-05-13T02:43:33-0400
Let g(x)=sqrt(x-2)=y. (1)

We need to find f(y). So, we substitute (1) in f(y)=y^2+3:

f(g(x))=(sqrt(x-2))^2+3=x-2+3=x+1.